OpenStudy (anonymous):

I multiply two consecutive numbers and get a product of 1122. What are the two numbers?

OpenStudy (anonymous):

First find the square root of 1122. The 2 numbers will surround the number you got as the root.

OpenStudy (unklerhaukus):

let the first number be \(n\) the other number is consecutive so it is \(n+1\) \[(n)\times(n+1)=1122\\n^2+n-1122=0\] apply the quadratic formula \[ax^2+bx+c=0\] \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] NB; there will be two solutions

mathslover (mathslover):

Hello @Xcresent , great to see you here on OS . \(\LARGE{\color{blue}{\mathsf{Welcome}} \space \color{orange}{\textbf{To}} \space \color{red}{\cal{OpenStudy}} \quad \ddot\smile}\)